Kingtaker435:

A closed container has 3.06 ⋅ 1023 atoms of a gas. Each atom of the gas weighs 1.67 ⋅ 10–24 grams. Which of the following shows and explains the approximate total mass, in grams, of all the atoms of the gas in the container?

Kingtaker435:

its math but it wont let me post

Mercury:

answer choices 0.47 grams, because (3.06 + 1.67) ⋅(1023 ⋅10–24) = 4.73 ⋅10–1 5.11 grams, because (3.06 ⋅1.67) ⋅(1023 ⋅10–24) = 5.1102 0.51 grams, because (3.06 ⋅1.67) ⋅(1023 ⋅10–24) = 5.1102 ⋅10–1 4.73 grams, because (3.06 + 1.67) ⋅(1023 ⋅10–24) = 4.73

Mercury:

multiply (# of atoms) * (mass per atom) to get total mass.